OpenStudy (deathdefire):

Algebra Question Help Please! Suppose a white dwarf star has a diameter of approximately 1.2519 × 10^4 km. Use the formula 4πr 2 to find the approximate surface area of the star.

Vocaloid (vocaloid):

hint: radius = (1/2) * diameter if diameter = 1.2519 × 10^4 km, then what is the radius?

OpenStudy (deathdefire):

6259.5

Vocaloid (vocaloid):

good now, if surface area = 4 * pi * (radius)^2 what is the surface are?

OpenStudy (deathdefire):

492,117,633.54 so the answer would be 4.92 * 10^8 km^2 right?

Vocaloid (vocaloid):

good

OpenStudy (deathdefire):

thanks!

OpenStudy (deathdefire):

could you help me with one more real quick? x^12y^-11/x^20y^-13